In the Hab:
A fan was added to the crew quarter air vent, increasing comfort in the morning. A single crew member still found it necessary to still sleep with the door open for air circulation.
The temperature in the Hab is stable and comfortable. The temperature was not adjusted since the start of our mission. The current inside temperature is 23C.
The caulking in the shower is in need of replacement. Request permission to strip the edges around the bottom and apply new caulking.
This morning’s cheese bread and this evening’s carrot cake projects were a complete success.
Trailer water tank: 5/40 in DG has left with the trailer.
Propane tank: 22%
Generator diesel: 70%
Yesterday’s final water use turned out to be 40 L/person. (1 shower, 1 sponge bath)
Today’s current waster use: 24 L/person. (1 sponge bath), this will increase throughout the evening.
Measures for reducing water consumption:
Grey water collection in the shower works well with sponge baths.
We have decided to wash dishes only twice a day.
The sink valves have been partially closed to reduce water flow, and a plastic diffuser is being used in the kitchen faucet for better rinsing.
All ATVs started. ATV 2 sputters. ATV red idle speed is low and stalls.
Filling the standing water on the south side of the Hab with dirt worked well. A final dirt load should solve this problem completely.
My previous fog mitigation system (magnets in a sock) has been improved upon with tightly sewn fabric around the magnets. This worked much better, and is no longer a choking hazard. Once the fabric is wet, the system looses efficiency. I will continue experimenting.
Awaiting instructions for a construction project.